The height of a triangle is four time the length of the base. Find length and the height of the triangle if the area is 18 square inches.
Answer by mananth(16949)

The height of a triangle is four time the length of the base. Find length and the height of the triangle if the area is 18 square inches.
let the base be x inches
height is 4x
Area of triangle = 1/2 Base * height
1/2 * x * 4x= 18
4x^2/2 = 18
4x^2=36
x^2= 9
x=3 inches . This is the base
The height will be 4*3 = 12 inches